On average, 3 traffic accidents per month intersection. What is the probability that in any given month at this intersection between 3 and 4 accidents occur?
1
Expert's answer
2020-11-17T06:07:58-0500

Let "X" be the random variable representing the number of accidents in a month: "X\sim Po(\lambda t)"


"P(X=x)=\dfrac{e^{-\lambda t}(\lambda t)^x}{x!}"

Average number of accidents per month, "\lambda=3, t=1."

Then "\lambda t=3(1)=3" and


"P(X=x)=\dfrac{e^{-3}(3)^x}{x!}"

"P(3\leq X\leq4)=P(X=3)+P(X=4)="

"=\dfrac{e^{-3}(3)^3}{3!}+\dfrac{e^{-3}(3)^4}{4!}=\dfrac{63e^{-3}}{8}\approx0.392073"

The probability that in any given month at this intersection between 3 and 4 accidents occur is "0.392073."